write each equation in standard form.
y+2=-3(x-1)

Answer by mananth(16946)

write each equation in standard form.
y+2=-3(x-1)
y+2=-3x+3
-2 -2
y+2-2=-3x+3-2
y=-3x+1
+3x +3x
y+3x=1
3x+y=1
...
m.ananth@hotmail.ca